Synchronization and Pipeline Design for a Multithreaded Massively Parallel Computer

Computation Structures Group Memo 343-1 March 1992

Shuichi Sakai

This report describes research done at the Laboratory for Computer Science of the Massachusetts Institute of Technology. Funding for the Laboratory is provided in part by the Advanced Research Projects Agency of the Department of Defense under the Office of Naval Research contract N00014-89-J-1988.